OpenStudy (anonymous):

A rectangular room on a blueprint is 2 inches wide and 5 inches long. If the scale of the blueprint to the actual room is 1:72, what is the width of the room?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

answer my question

OpenStudy (anonymous):

Area of room = length * breadth =5*2=10

OpenStudy (anonymous):

that is not the answer

OpenStudy (anonymous):

i have not solve. it its just starting

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oops i did wrong

OpenStudy (anonymous):

now see , 2 / x = 1 / 72 x = 144 144 inches = 12 feet The width of the room is 12 f

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

welcum